# SSD: A State-based Stealthy Backdoor Attack For Navigation System in UAV Route Planning

### v1

@@ Abstract
Unmanned aerial vehicles (UAVs) are increasingly employed to perform high-risk tasks that require minimal human intervention. However, UAVs face escalating cybersecurity threats, particularly from GNSS spoofing attacks. While previous studies have extensively investigated the impacts of GNSS spoofing on UAVs, few have focused on its effects on specific tasks. Moreover, the influence of UAV motion states on the assessment of network security risks is often overlooked. To address these gaps, we first provide a detailed evaluation of how motion states affect the effectiveness of network attacks. We demonstrate that nonlinear motion states not only enhance the effectiveness of position spoofing in GNSS spoofing attacks but also reduce the probability of speed-related attack detection. Building upon this, we propose a state-triggered backdoor attack method (SSD) to deceive GNSS systems and assess its risk to trajectory planning tasks. Extensive validation of SSD’s effectiveness and stealthiness is conducted. Experimental results show that, with appropriately tuned hyperparameters, SSD significantly increases positioning errors and the risk of task failure, while maintaining high stealthy rates across three state-of-the-art detectors.
1 Introduction
@@
Unmanned Aerial Vehicles (UAVs) are revolutionizing our understanding of low-altitude flight patterns. Currently, UAVs are widely used in various military and civilian fields, such as courier delivery, agricultural plant protection, power patrol, firefighting, rescue, and battlefield reconnaissance. These increasingly complex application scenarios have necessitated stringent requirements for the autonomous flight capabilities of UAVs. As UAVs operate autonomously and execute their tasks, the system must precisely determine its global position at a centimeter level. This localization capability is critical to the safety of the flight and its ability to fulfill its mission, as positioning errors can directly cause the flight to deviate from course or fail to perform its mission.
@@
The Integrated Navigation System (INS) serves as the cornerstone of UAVs, enabling precise positioning. It accomplishes accurate position estimates by integrating data from various sensors. Integrating Inertial Measurement Units (IMUs) and Global Navigation Satellite Systems (GNSS) forms the fundamental and core navigation system in INS. Building on this foundation, researchers have incorporated additional sensors, such as cameras, lidar, and radar to further improve positioning accuracy. However, direct reliance on sensor data and communication channels’ noise makes INS vulnerable to cyber-attacks [1]. Research has revealed that adversaries can attack INS by using adversarial examples or wireless signal injection to spoof sensors. Notably, GNSS spoofing is a particularly prevalent threat since it forms the foundation of INS. The attacker can leverage low-cost devices to manipulate the position and velocity measurement captured by GNSS.
@@
Previous attacks can be classified into the following categories:(1) Direct Attacks[2] : The adversarial directly injects a false signal to the GNSS sensor. (2) Stealthy Attacks [3, 4]: The adversarial takes the detector and corresponding threshold as the constraint and computes an optimization-based payload to bypass the detectors.
Figure 1: Route Planing
@@
However, these attacks have the following limitations.(1) Easily detectable Sensor data fluctuations during UAV attacks are typically significant, prompting residual-based detection methods. However, our experiments show that these methods can almost always detect direct attacks. Moreover, during UAV maneuvers, such as waypoint course changes or formation shifts, the increased fluctuations make attacks even more easily identifiable. (2) Computation efficiency While stealthy attacks can bypass detectors through constrained optimization models, they typically require extensive matrix operations to solve high-dimensional optimization problems and derive the optimal attack payload in time. These methods introduce significant computational delays in environments with limited resources, especially affecting the attack’s real-time performance. (3) Inadequate analysis of dynamic vulnerabilities Research [5] has shown that navigation algorithms in autonomous vehicles are vulnerable to uncertainty during specific periods. UAVs, operating with six degrees of freedom, experience frequent motion changes that affect system stability, especially during GNSS spoofing. Despite this, few studies evaluate how these dynamic motion states influence attack effectiveness. (4) Incomplete assessment methodology Prior studies often focus on immediate attack outcomes, such as crashes or path deviations, but fail to assess how attacks affect UAV mission performance and overall efficacy. This leaves a gap in understanding the broader impacts of such attacks on UAV operations.
@@
To overcome these limitations, and gain a deep understanding of the INS vulnerability posed by GNSS attacks, we first provided a detailed interpretable analysis of the relationship between motion states and GNSS spoofing attacks. Specifically, we assess the effects of linear and nonlinear motion states on attack effectiveness. Our findings reveal that changes in motion states amplify the effectiveness of positional attacks and increase the stealthiness of velocity attacks. We then proposed SSD, a novel stated-based stealthy backdoor attack for GNSS. Backdoor attacks [6] are a common threat in deep neural networks, where an adversary implants a latent backdoor that remains inactive under normal conditions but is triggered by specific inputs or scenarios, leading to incorrect model predictions. Inspired by this concept, we first design backdoors for GNSS by using motion state changes as a trigger to initiate staged velocity and positional attacks. In contrast to previous work [4], this attack mode doesn’t need prior knowledge for the detectors like stealthy attacks. Besides, it eliminates the need for complex computations. Instead, it relies on the reasonable configuration of parameters and simple function calculations to strike an optimal balance between effectiveness and stealthiness. This makes SSD highly valuable for engineering applications. Lastly, we selected three representative task trajectories to assess the effectiveness of SSD. Its performance was compared against existing attack methods. The experimental results demonstrate that SSD maintains detection variables consistently within the threshold range in three classical detectors. Furthermore, we introduced evaluation metrics to measure the attack’s impact on mission success rates and effectiveness. The experiments also reveal that SSD significantly increases the localization error, thereby effectively disrupting mission completion.
@@
  • We present an interpretable mathematical security study of how motion states influence attack outcomes and demonstrate that UAVs are more vulnerable during maneuvers than in uniform linear flights. We further experimentally proof it.
  • We design SSD, a novel stated-based backdoor attack that utilizes motion state as a trigger to spoof GNSS data. It can execute velocity and positional attacks in stages to simultaneously and covertly attack both two parameters.
@@
2 Related Works and Background
2.1 UAV Route Planning and Integrated Navigation System
@@
The UAV route planning involves designing a feasible route from the start point to the destination while meeting all constraints and performance requirements [7]. Effective route planning is crucial for UAVs to complete their missions, while accurate position estimation is the basis for achieving efficient and accurate route planning. Since UAVs often operate in dynamic and complex environments, they rely on multi-sensor fusion to enhance their ability to perceive and process information. This approach integrates data from various sensors with different modalities and attributes, increasing redundancy and improving reliability in challenging conditions. The integrated navigation system of GNSS and IMU is a typical representative example. Its fusion strategy uses the GNSS data for quantitative updating, IMU data for state prediction, and optimal estimation framework to achieve accurate positioning in the global coordinate system [8, 9]. In GPS-denied environments, simultaneous localization and mapping (SLAM) that rely on Camera [10, 11] and Lidar [12, 13] are considered more reliable solutions. However, a single sensor alone cannot fully meet the demands for positioning accuracy and response speed. Therefore, combining these sensors with an inertial measurement unit to create IMU/Camera [14, 15] and IMU/Lidar [16] fusion form enables more precise position estimation and improved performance in dynamic environments. This paper focuses on the security analysis of IMU/GNSS INS since it plays a central role in UAVs. It is of generic meaning to study its security.
2.2 GNSS Spoofing attack
@@
IMUs are more resilient to manipulation in real-world scenarios. For IMU/GNSS INS, GNSS spoofing attacks on UAVs are serious safety and security threats. In such attacks, the adversary transmits false location coordinates to the GNSS receiver, causing it to process the fraudulent data as legitimate, thereby concealing the UAV’s true location. As a result, unknowingly accepting these false inputs, the navigation system calculates incorrect position information. Specifically, GNSS spoofing includes direct attacks and stealthy attacks. The direct attacks [2] have the following categories: (1) Biased Signal Attack : These attacks involve adding a bias to the GNSS sensor signals, typically following a uniform distribution. (2) Multiplicative Attacks: In these attacks, the GNSS signals are multiplied by a constant factor, effectively scaling the original signal values. (3) Replacement Attacks: These attacks involve directly replacing the GNSS signals with false or manipulated data.
The stealthy attacks [3, 4] are diverse, with attackers often aiming to maximize navigation residuals to determine the optimal attack sequence. The design of these attacks is largely influenced by the detection mechanism and its threshold. All these attacks can lead to disastrous consequences, such as the UAV crashing into obstacles or falling into the malicious attackers’ control [17, 18, 19].
@@
To mitigate this threat, robust countermeasures have been implemented to safeguard against GNSS spoofing and ensure the integrity of UAV INS. One important method is multi-sensor fusion. As described in section 2.1, it provides more accurate estimates for sensing and localization, for example, and enhances the data’s trustworthiness, providing greater redundancy for detection and defense in the event of spoofing attacks. For example, Shen et al. [5] demonstrated that the effectiveness of constant offset GNSS spoofing attacks is greatly reduced in GNSS/INS/LiDAR fused navigation system in autonomous driving. However, such fusion multi-sensor fusion strategy also faces a period of vulnerability and does not defend against attacks constructed GNSS spoof attack against the uncertainty that exists in the fusion algorithm itself. In this paper, we demonstrate a similar phenomenon in UAVs, where changes in motion states significantly amplify the uncertainty of the INS, making it more vulnerable to GNSS attacks. Therefore, we provide an in-depth analysis of the uncertainty and vulnerability and exploit it to design SSD.
2.3 Threat Model
@@
Attack Goal As shown in Figure 2, the adversarial aims to make the drone deviate significantly from its pre-planned route without triggering the stealth detection threshold. This objective can be formalized as the following optimization problem:
@@
where $\mathbf{D^{t}_{i}}$ and $\mathbf{D^{a}_{i}}$ denote the normal trajectory and the attacked trajectory, respectively. $\delta$ is the attack payload. $\chi_{k}$ is the in detection statistics and $\tau$ is the threshold of detectors
@@
Attack Senario As shown in Figure 2, an attacker can launch an attack in two ways: ① The attacker can use a UAV to fly alongside the victim’s UAV. He can transmit legitimate GNSS signals completely using wireless attack devices such as software-defined radios (SDR). ② The attacker could also inject specific backdoors [20] or viruses [21] into the UAV by supply chain attacks. that could be used to monitor its dynamics and induce a false GNSS position [22].
Attacker’s Capability 1)The attackers need white-box access to obtain the victim’s navigation algorithms and corresponding parameters. They can get this knowledge through open-source channels since most UAVs use standardized open-source navigation algorithms [23, 24]. Also the adversarial can use reverse engineering to access the victim’s knowledge. 2) The attacker can obtain the victim’s motion state such as position and velocity. This can be achieved by monitoring UAVs using an additional GPS module or auxiliary object detection and tracking devices. 3) As IMU data is less likely to be accessed and used, an attacker can only modify the position and velocity since GNSS measurement only provide position and velocity to the UAV.
3 Preliminary
3.1 IMU/GNSS Integrated Navigation System
In UAVs, IMU and GNSS are often combined for highly accurate and robust navigation and positioning. IMU provides data from accelerometers and gyroscopes, which measure the acceleration and angular velocity of the carrier. At the same time, GNSS determines the UAV’s position, velocity, and time information by receiving satellite signals. This type of navigation system is known as a combined IMU/GNSS navigation system.
@@ 3.1 IMU/GNSS Integrated Navigation System
TABLE I: Axis and meaning of EKF
Axis Label Meaning
1 $\mathbf{q_{0}}$ Orientation quaternion.
2 $\mathbf{q_{1}}$
3 $\mathbf{q_{2}}$
4 $\mathbf{q_{3}}$
5 $\mathbf{P_{N}}$ UAV Position in local NED coordinate system.
6 $\mathbf{P_{E}}$
7 $\mathbf{P_{D}}$
8 $\mathbf{v_{N}}$ UAV Velocity in local NED coordinate system.
9 $\mathbf{v_{E}}$
10 $\mathbf{v_{D}}$
11 $\mathbf{\Delta\theta bias_{X}}$ Bias in integrated gyroscope reading.
12 $\Delta\theta bias_{Y}$
13 $\Delta\theta bias_{Z}$
14 $\Delta vbias_{X}$ Bias in integrated accelerometer reading.
15 $\Delta vbias_{Y}$
16 $\Delta vbias_{Z}$
17 $geomagneticFieldVector_{N}$ Estimate of geomagnetic field vector at the reference location.
18 $geomagneticFieldVector_{E}$
19 $geomagneticFieldVector_{D}$
20 $magbias_{X}$ Bias in the magnetometer readings.
21 $magbias_{Y}$
22 $magbias_{Z}$
@@
Due to hardware arithmetic limitations, INS often handle nonlinear functions by truncating their Taylor expansions with first-order linearization and neglecting the higher-order terms. This approach transforms the nonlinear problem into a linear one, as exemplified by Equation 4 and 5:
@@
INS defines the state transfer matrix $F_{K}$ and Jacobi matrix of the measurement $H_{k}$ respectively:

### v2

# SSD: A State-based Stealthy Backdoor Attack For IMU/GNSS Navigation System in UAV Route Planning

@@ Abstract
Unmanned aerial vehicles (UAVs) are increasingly employed to perform high-risk tasks that require minimal human intervention. However, they face escalating cybersecurity threats, particularly from GNSS spoofing attacks. While previous studies have extensively investigated the impacts of GNSS spoofing on UAVs, few have focused on its effects on specific tasks. Moreover, the influence of UAV motion states on the assessment of cybersecurity risks is often overlooked. To address these gaps, we first provide a detailed evaluation of how motion states affect the effectiveness of network attacks. We demonstrate that nonlinear motion states not only enhance the effectiveness of position spoofing in GNSS spoofing attacks but also reduce the probability of detecting speed-related attacks. Building upon this, we propose a state-triggered backdoor attack method (SSD) to deceive GNSS systems and assess its risk to trajectory planning tasks. Extensive validation of SSD’s effectiveness and stealthiness is conducted. Experimental results show that, with appropriately tuned hyperparameters, SSD significantly increases positioning errors and the risk of task failure, while maintaining high stealthy rates across three state-of-the-art detectors.
Index Terms:
Unmanned aerial vehicles, Cyber security, Backdoor attacks, GNSS spoofing.
I Introduction
@@
Unmanned Aerial Vehicles (UAVs) are revolutionizing our understanding of low-altitude flight patterns. Currently, UAVs are widely used in various military and civilian fields, such as courier delivery, agricultural plant protection, power patrol, firefighting, rescue, and battlefield reconnaissance. These increasingly complex application scenarios have necessitated stringent requirements for the autonomous flight capabilities of UAVs. As UAVs operate autonomously to execute their missions, the system must precisely determine its global position at a centimeter level. As illustrated in Fig. 1,
Figure 1: The Role of Localization and Route Planing in UAV Autonomous Flight
the localization capability is critical to route planning. It ensures the safety of flight and the ability to fulfill its mission, as positioning errors can directly cause the flight to deviate from course or fail to perform its mission.
@@
The Integrated Navigation System (INS) serves as the cornerstone of UAVs, enabling precise positioning. It accomplishes accurate position estimates by integrating data from various sensors. Integrating Inertial Measurement Units (IMUs) and Global Navigation Satellite Systems (GNSS) forms the fundamental and core navigation system in INS. Building on this foundation, researchers usually incorporated additional sensors, such as cameras, lidar, and radar to further improve positioning accuracy in different scenarios or platforms. However, direct reliance on sensor data and communication channels’ noise makes INS vulnerable to cyber-attacks [1, 2]. Research has revealed that adversaries can attack INS by using adversarial examples [3] or wireless signal injection [4] to spoof sensors. Notably, GNSS is a particularly prevalent threat since it forms the foundation of INS. The attacker can leverage low-cost devices to manipulate the position and velocity measurement captured by GNSS.
@@
Previous attacks can be classified into the following categories:(1) Direct Attacks[5] : The adversary directly injects a false signal into the GNSS sensor. (2) Stealthy Attacks [6, 7]: The adversary takes the detector and corresponding threshold as the constraint and computes an optimization-based payload to bypass the detectors.
@@
However, these attacks have the following limitations.(1) Easily detectable Sensor data fluctuations during UAV attacks are typically significant [8], prompting residual-based detection methods. However, our experiments show these detectors can almost always detect direct attacks. Moreover, during UAV maneuvers, such as waypoint course changes or formation shifts, the increased fluctuations make attacks even more easily identifiable. (2) Computation efficiency While stealthy attacks can bypass detectors through constrained optimization models, they typically require extensive matrix operations to solve high-dimensional optimization problems and derive the optimal attack payload in time. These methods introduce significant computational delays in environments with limited resources, especially affecting the attack’s real-time performance. (3) Inadequate analysis of dynamic vulnerabilities Research [9] has shown that navigation algorithms in autonomous vehicles are vulnerable to uncertainty during specific periods. UAVs, operating with six degrees of freedom, experience frequent motion changes that affect system stability, especially during GNSS spoofing. Despite this, few studies evaluate how these dynamic motion states influence attack effectiveness. (4) Incomplete assessment methodology Prior studies [10, 11] often focus on immediate attack outcomes, such as crashes or path deviations, but fail to assess how attacks affect UAV mission performance and overall efficacy. This leaves a gap in understanding the broader impacts of such attacks on UAV operations.
@@
To overcome these limitations, and gain a deep understanding of the INS vulnerability posed by GNSS attacks, we first provided a detailed interpretable analysis of the relationship between motion states and GNSS spoofing attacks. Specifically, we assess the effects of linear and nonlinear motion states on attack effectiveness. Our findings reveal that changes in motion states amplify the effectiveness of positional attacks and increase the stealthiness of velocity attacks. We then proposed SSD, a novel stated-based stealthy backdoor attack for GNSS. Backdoor attacks [12, 13] are a common threat in deep neural networks, where an adversary implants a latent backdoor that remains inactive under normal conditions but is triggered by specific inputs or scenarios, leading to incorrect model predictions. Inspired by this concept, we design backdoors for GNSS by using motion state changes as a trigger to initiate staged velocity and positional attacks. In contrast to the stealthy attacks [6, 7], this attack mode doesn’t need prior knowledge for the detectors and eliminates the need for complex computations. Instead, it is a direct attack that leverages carefully configured parameters and straightforward function calculations to achieve an optimal balance between effectiveness and stealthiness. This makes SSD highly valuable for engineering applications. Lastly, we selected three representative mission trajectories to assess the effectiveness of SSD. Its performance was compared against existing attack methods. The experimental results demonstrate that SSD maintains detection variables consistently within the threshold range in three classical detectors. Furthermore, we introduced evaluation metrics to measure the attack’s impact on mission success rates and effectiveness. The experiments also reveal that SSD significantly increases the localization error, thereby effectively disrupting mission completion.
@@
  • We present an interpretable mathematical security study of how motion states influence attack outcomes and demonstrate that UAVs are more vulnerable during maneuvers than in uniform linear flights. We further experimentally prove it.
  • We design SSD, a novel stated-based backdoor attack that utilizes motion state as a trigger to spoof GNSS data. It can execute velocity and positional attacks in stages to simultaneously and covertly attack both states.
@@
II Related Works and Background
II-A UAV Route Planning and Integrated Navigation System
@@
UAV route planning involves designing a feasible route from the start point to the destination while meeting all constraints and performance requirements [14]. Effective route planning is crucial for UAVs to complete their missions, and it depends on accurate position estimation. Since UAVs often operate in dynamic and complex environments, they rely on multi-sensor fusion to enhance their ability to perceive the environment. This approach integrates data from various sensors with different modalities and attributes, increasing redundancy and improving reliability in challenging conditions. The integrated navigation system of GNSS and IMU is a typical representative example. Its fusion strategy uses the GNSS data for quantitative updating, IMU data for state prediction, and an optimal estimation framework to achieve accurate positioning in the global coordinate system [15, 16]. In GNSS-denied environments, simultaneous localization and mapping (SLAM) that rely on camera [17, 18] and lidar [19, 20] are considered more reliable solutions. However, a single sensor alone cannot fully meet the demands for positioning accuracy and response speed. Therefore, combining these sensors with an inertial measurement unit to create IMU/Camera [21, 22] and IMU/Lidar [23] fusion form enables more precise position estimation and improved performance in dynamic environments. This paper focuses on the security analysis of IMU/GNSS INS since it plays a central role in UAVs. It is of generic meaning to study its security.
II-B GNSS Spoofing attack
@@
Since IMUs are more difficult to manipulate in real-world scenarios, we only discuss GNSS spoofing attacks for IMU/GNSS INS on UAVs. In such attacks, the adversary transmits false location coordinates to the GNSS receiver, thereby concealing the UAV’s true location. As a result, unknowingly accepting these false inputs, the navigation system calculates incorrect position information. Specifically, GNSS spoofing includes direct attacks and stealthy attacks. The direct attacks [5] can be classified into these categories. (1) Biased Signal Attack : These attacks involve adding a bias to the GNSS sensor signals, typically following a uniform distribution. (2) Multiplicative Attacks: In these attacks, the GNSS signals are multiplied by a constant factor, effectively scaling the original signal values. (3) Replacement Attacks: These attacks involve directly replacing the GNSS signals with false or manipulated data. Direct attacks are easy to implement and may lead to disastrous consequences, such as the UAV crashing into obstacles [10, 11]. However, most detectors can detect and respond in time. The stealthy attacks [6, 7] are diverse, with attackers often aiming to maximize navigation residuals to determine the optimal attack sequence. The design of these attacks may cause the UAV to fall into the malicious attackers’ control [24]. However, it is largely influenced by the detection mechanism and often needs complex computing. SSD integrates the strengths of both approaches. It achieves the same effectiveness as stealthy attacks while requiring low computational resources.
@@
To mitigate this threat, robust countermeasures, such as software analysis [25], cryptography-based authentication [26] and machine learning-based detections [27, 28], have been implemented to safeguard against GNSS spoofing and ensure the integrity of UAV INS. One important method is multi-sensor fusion [29]. As described in section II-A, it not only provides more accurate estimates for perception and localization. but also enhances the data’s trustworthiness, providing greater redundancy for detection and defense in the event of spoofing attacks. For example, Shen et al. [9] demonstrated that the effectiveness of constant offset GNSS spoofing attacks is greatly reduced in GNSS/INS/LiDAR fused navigation systems in autonomous driving. However, such multi-sensor fusion strategies also face a period of vulnerability and can not defend against constructed GNSS spoofing against the uncertainty that exists in the fusion algorithm itself. In this paper, we demonstrate a similar phenomenon in UAVs, where changes in motion states significantly amplify the uncertainty of the INS, making it more vulnerable to GNSS attacks. Therefore, we provide an in-depth analysis of the uncertainty and vulnerability and exploit it to design SSD.
II-C Threat Model
@@
Attack Goal As shown in Figure 2, the adversary aims to make the drone deviate significantly from its pre-planned route without triggering the stealth detection threshold. This objective can be formalized as the following optimization problem:
@@
where $\mathbf{D^{t}_{i}}$ and $\mathbf{D^{a}_{i}}$ denote the normal trajectory and the attacked trajectory, respectively. $\delta$ is the attack payload. $\chi_{k}$ is the in detection statistics and $\tau$ is the threshold of detectors.
@@
Attack Scenario As shown in Fig. 2, an attacker can launch an attack in two ways: ① The attacker can use a UAV to fly alongside the victim’s UAV. He can transmit legitimate GNSS signals completely using wireless attack devices such as software-defined radios (SDR). ② The attacker could also inject specific backdoors [30] or viruses [31] into the UAV by supply chain attacks, which could be used to monitor its dynamics and induce a false GNSS position [32].
Attacker’s Capability 1) The attackers need white-box access to obtain the victim’s navigation algorithms and corresponding parameters. They can get this knowledge through open-source channels since most UAVs use standardized open-source navigation algorithms [33, 34]. Also, the adversarial can use reverse engineering to access the victim’s knowledge. 2) The attacker can obtain the victim’s motion state, such as position and velocity. This can be achieved by monitoring UAVs using an additional GPS module or auxiliary object detection and tracking devices. 3) As IMU data is less likely to be accessed and used, an attacker can only modify the position and velocity since GNSS measurements only provide position and velocity to the UAV.
III Preliminary
III-A IMU/GNSS Integrated Navigation System
In UAVs, IMU and GNSS are often combined for highly accurate and robust navigation and positioning. IMU provides data from accelerometers and gyroscopes, which measure the acceleration and angular velocity of the UAVs. At the same time, GNSS determines the UAV’s position, velocity, and time information by receiving satellite signals. This type of navigation system is known as a combined IMU/GNSS navigation system.
@@ III-A IMU/GNSS Integrated Navigation System
TABLE I: Element and Meaning of EKF Vector
Element Label Meaning
$\hat{x}_{1}$ $\mathbf{q_{0}}$ Orientation quaternion.
$\hat{x}_{2}$ $\mathbf{q_{1}}$
$\hat{x}_{3}$ $\mathbf{q_{2}}$
$\hat{x}_{4}$ $\mathbf{q_{3}}$
$\hat{x}_{5}$ $\mathbf{P_{N}}$ UAV Position in local NED coordinate system.
$\hat{x}_{6}$ $\mathbf{P_{E}}$
$\hat{x}_{7}$ $\mathbf{P_{D}}$
$\hat{x}_{8}$ $\mathbf{V_{N}}$ UAV Velocity in local NED coordinate system.
$\hat{x}_{9}$ $\mathbf{V_{E}}$
$\hat{x}_{10}$ $\mathbf{V_{D}}$
$\hat{x}_{11}$ $\mathbf{\Delta\theta bias_{X}}$ Bias in integrated gyroscope reading.
$\hat{x}_{12}$ $\mathbf{\Delta\theta bias_{Y}}$
$\hat{x}_{13}$ $\mathbf{\Delta\theta bias_{Z}}$
$\hat{x}_{14}$ $\mathbf{\Delta vbias_{X}}$ Bias in integrated accelerometer reading.
$\hat{x}_{15}$ $\Delta vbias_{Y}$
$\hat{x}_{16}$ $\mathbf{\Delta vbias_{Z}}$
$\hat{x}_{17}$ $\mathbf{geomagneticField_{N}}$ Estimate of geomagnetic field vector at the reference location.
$\hat{x}_{18}$ $\mathbf{geomagneticField_{E}}$
$\hat{x}_{19}$ $\mathbf{geomagneticField_{D}}$
$\hat{x}_{20}$ $\mathbf{magbias_{X}}$ Bias in the magnetometer readings.
$\hat{x}_{21}$ $\mathbf{magbias_{Y}}$
$\hat{x}_{22}$ $\mathbf{magbias_{Z}}$
@@
Due to hardware arithmetic limitations, INS often handles nonlinear functions by truncating their Taylor expansions with first-order linearization and neglecting the higher-order terms. This approach transforms the nonlinear problem into a linear one, as exemplified by Eq. 4 and 5:
@@
INS defines the state transfer matrix $F_{K}$ and the Jacobi matrix of the measurement $H_{k}$ respectively: